### v1

# EPN: An Ego Vehicle Planning-Informed Network for Target Trajectory Prediction

Abstract
Trajectory prediction plays a crucial role in improving the safety and reliability of autonomous vehicles, serving as an intermediate link between perception and planning. However, due to the highly dynamic and multimodal nature of the task, accurately predicting the future trajectory of a target vehicle remains a significant challenge. To address these challenges, we propose an Ego vehicle Planning-informed Network (EPN) for multimodal trajectory prediction. Current trajectory prediction methods typically use the historical trajectory and vehicle attributes as inputs, focusing primarily on how historical information influences the future trajectory of the target vehicle. In real-world driving scenarios, however, the future trajectory of a vehicle is influenced not only by its own historical data but also by the behavior of other vehicles on the road. To address this, we incorporate the future planned trajectory of the ego vehicle as an additional input to simulate the mutual influence between the ego vehicle’s planned trajectory and the predicted trajectory of the target vehicle. Furthermore, to tackle the challenges of intention ambiguity and large prediction errors often encountered in methods based on driving intentions, we propose a target’s endpoint prediction module. This module first predicts the possible endpoints of the target vehicle, then refines these predictions through a correction mechanism, and finally generates a complete multimodal predicted trajectory based on the corrected endpoints. Experimental results demonstrate that, compared to other trajectory prediction methods, EPN achieves an average reduction of 34.9%, 30.7%, and 30.4% in RMSE, ADE, and FDE evaluation metrics on the NGSIM dataset, and an average reduction of 64.6%, 64.5%, and 64.3% in RMSE, ADE, and FDE on the HighD dataset. These results highlight the strong performance of EPN in trajectory prediction. The code is available at https://github.com/whutpsq/EPN.
Index Terms:
Autonomous driving, trajectory prediction, ego vehicle planning, endpoint correction, LSTM, CVAE.
I Introduction
With the rapid development of autonomous driving technology, the safety of autonomous vehicles has garnered increasing attention. Accurately predicting the future trajectory of surrounding vehicles is crucial for the safe operation of autonomous vehicles[2]. As an upstream module of planning, trajectory prediction plays a key role in supporting planning tasks and is one of the critical components in enhancing the safety of autonomous driving systems. However, due to the high dynamic and multimodal nature of trajectory prediction tasks, accurately forecasting the future trajectory of a target vehicle remains a significant challenge. High dynamism refers to the influence of other traffic participants on the target vehicle, which may cause the target vehicle’s trajectory to change unexpectedly. Multimodality refers to the possibility of multiple plausible future trajectories for a given historical trajectory. These characteristics make trajectory prediction a complex and difficult task.
In the early stages of trajectory prediction technology, methods based on physical models were commonly used. Dynamic models made short-term predictions by describing the vehicle’s dynamic characteristics[3][4]. However, these models are often complex, and their performance improvement in trajectory prediction is limited despite the increased model complexity. Kinematic models, on the other hand, are simpler, primarily using kinematic attributes such as vehicle speed, acceleration, and steering angle[5][6]. However, these models are only suitable for relatively simple scenarios. These traditional methods primarily focus on the state of the target vehicle and fail to account for the influence of other factors, such as surrounding vehicles and environmental conditions. The Kalman filter method incorporates noise consideration within the model construction and predicts vehicle trajectories through Gaussian distribution modeling.[7][8]. Compared to the Kalman filter, which is limited to linear scenarios, the Monte Carlo method offers advantages in handling complex nonlinear and multimodal scenarios[9]. However, practical applications require a balance between efficiency and accuracy.
As technology has advanced, classic machine learning methods have been increasingly applied to vehicle trajectory prediction. Unlike physics-based models, machine learning methods are data-driven approaches. Notable methods include Gaussian Processes (GP), Support Vector Machines (SVM), Hidden Markov Models (HMM), and Dynamic Bayesian Networks (DBN). GP use historical trajectory data to learn the potential distribution of future trajectories[10]. SVM classify data by finding an optimal hyperplane, helping to determine driving intentions such as left turns, straight movements, and right turns[11]. HMM effectively simulate temporal changes and uncertainties in driving behavior by modeling vehicle driving patterns as hidden states, but they are limited to systems with discrete states[12][13]. DBN are well-suited for handling complex multidimensional states and long-term dependencies, but they can suffer from significant errors when converting recognized intentions into accurate trajectories[14][15].
With the rapid development of deep learning, neural network models have made significant advancements in vehicle trajectory prediction, encompassing Recurrent Neural Network (RNN), Convolutional Neural Network (CNN), Graph Neural Network (GNN), Generative Adversarial Network (GAN) and Conditional Variational Auto Encoder (CVAE)[2][16]. Researchers often leverage one or more neural network architectures to achieve better prediction accuracy.
Owing to their extensive parameterization, neural networks can effectively model not only vehicle dynamics and attributes but also complex interactions among traffic participants and road environment features.
Deo et al.[17] proposed a convolutional social pooling module based on [18], which extracts the interaction relationships between vehicles at a deeper level. Ding et al.[19] introduced a vehicle behavior interaction network using recurrent neural networks to model interactions based on the hidden states of vehicles. Xin et al.[20] developed a trajectory prediction method utilizing two LSTMs, which first identifies the driver’s intention and then predicts the future trajectory of the vehicle based on that intention. Gupta et al.[21] applied generative adversarial networks to predict the trajectories of target vehicles. Wang et al.[22] proposed a social pooling method inspired by wave superposition to dynamically model both local and global interactions.
Incorporating prior knowledge significantly enhances trajectory prediction accuracy. For instance, high-definition maps have been widely used to improve predictions. Casas et al.[23] jointly tackled vehicle detection and trajectory prediction by processing raw nt clouds and gridded high-definition maps. Gao et al.[24] introduced vectorized encoding for high-definition maps and traffic participants, reducing information loss typically caused by rasterization, lowering computational complexity, and improving prediction accuracy. Subsequent studies, such as [25] -[28], extensively utilized vectorized encoding methods.
Beyond fully utilizing available information, researchers also aim to improve prediction accuracy by incorporating additional data. Song et al.[29] were among the first to consider how different ego vehicle plans impact the future trajectory of a target vehicle. They used an LSTM encoder to model the ego vehicle’s planned trajectory and extracted interaction features through a convolutional social pooling module. Their experiments demonstrated that incorporating ego vehicle planning information significantly improves model prediction accuracy. Guo et al.[30] further enhanced model performance by using a spatiotemporal dual-attention mechanism to extract spatiotemporal features while incorporating autonomous vehicle planning, achieving higher trajectory prediction accuracy. Sheng et al.[31] further reduced uncertainty in the target vehicle’s behavior by explicitly encoding ego vehicle’s planned trajectory into a graph convolutional network.
The characteristics of input information significantly influence trajectory prediction models, while the output format of these models is also crucial in determining their performance. Multimodal trajectory prediction typically employs two approaches. The first is driving intention-based prediction, where driver behaviors such as left turns, right turns, straight movements, acceleration, and deceleration are classified, and trajectories are predicted for each intention, as seen in studies like [29], [30], [32] and [33]. The second is endpoint-based prediction, such as [26], [34] and [35], which first predict the target vehicle’s potential endpoint and subsequently generate the complete trajectory based on the predicted endpoint.
In this paper, we propose a multimodal trajectory prediction model based on endpoint correction and ego vehicle planning. This model takes as input the historical trajectories and state information of all vehicles in the scene, along with the planned trajectory of the ego vehicle, and outputs multimodal predictions for the future trajectory of the target vehicle. We intentionally did not incorporate high-definition maps due to their high production and updating costs. With the current trend in intelligent driving solutions moving towards lightweight maps, trajectory prediction methods that do not rely on high-definition maps still hold significant research value. Our contributions can be summarized as follows:
  1. 1.
    We propose a multimodal trajectory prediction model based on ego vehicle planning, significantly improving the prediction performance of mapless trajectory prediction methods.
  2. 2.
    We introduce a feature fusion encoding module that first extracts temporal features from the input information using an LSTM encoder, then models interactions between various traffic participants with a convolutional social pooling network to capture spatial features of the scene. The module outputs a unified feature vector that integrates both temporal and spatial characteristics.
  3. 3.
    We propose a target’s endpoint prediction module that first predicts the possible endpoints of the target vehicle using a CVAE, then refines the preliminary predicted endpoints coordinates with a correction mechanism, and finally uses the corrected endpoints to predict the complete trajectory of the target vehicle.
The remainder of the paper is organized as follows: Section II presents the problem formulation, Section III describes the model architecture, Section IV details the experimental setup and results, and Section V concludes the paper with a summary.
II PROBLEM FORMULATION
As illustrated in Figure 1, vehicles in the driving scenario are classified into three types: the red vehicle represents the ego vehicle, blue represents the target vehicle, and gray represents adjacent vehicles. The area of interest is divided into a grid of 200 by 35 feet, centered on the ego vehicle, denoted as $A_{tar}$, which corresponds to the region within the red dashed line. Similarly, the area surrounding the target vehicle, referred to as $A_{ner}$, is shown within the blue dashed line in the figure.
Figure 1: Vehicle Classification and Region of Interest Division in A Driving Scenario.
II-A Model Input
The model’s input consists of the historical driving state information of all vehicles in the scenario, represented as:
${X^{i}}=\{s_{t-{T_{h}}+1}^{i},s_{t-{T_{h}}+2}^{i},...,s_{t}^{i}\}$ (1)
where $i$ denotes the vehicle index, $T_{h}$ represents the historical time steps used as input, and $s_{t}^{i}$ is the driving state of the $i$-th vehicle at time step $t$, defined as:
$s_{t}^{i}=(x_{t}^{i},y_{t}^{i},v_{t}^{i},a_{t}^{i})$ (2)
where $x_{t}^{i},y_{t}^{i},v_{t}^{i},$and $a_{t}^{i}$ represent the horizontal position, vertical position, velocity, and acceleration of vehicle $i$ at time step $t$, respectively.
Additionally, the model input includes the future planned trajectory of the ego vehicle, given by:
$P=\{{p_{t+1}},{p_{t+2}},...,{p_{t+{T_{f}}}}\}$ (3)
where $T_{f}$ denotes the future prediction horizon, and $p_{t+T_{f}}$ is the planned position of the ego vehicle at time step $(t+T_{f})$, defined as:
${p_{t+T_{f}}}=({x_{t+T_{f}}},{y_{t+T_{f}}})$ (4)
where $x_{t+T_{f}}$,$y_{t+T_{f}}$ are the horizontal and vertical coordinates of the ego vehicle at time $(t+T_{f})$, respectively.
II-B Model Output
The model’s output is the predicted position sequence of the target vehicle, represented as:
$\widehat{Y}=\{\widehat{f}_{t+1}^{i},\widehat{f}_{t+2}^{i},...,\widehat{f}_{t+{T_{f}}}^{i}\}$ (5)
where $\widehat{f}_{t+{T_{f}}}^{i}$ denotes the predicted trajectory of the $i$-th target vehicle at time step $(t+T_{f})$. This predicted trajectory can be further expressed as:
$\widehat{f}_{t+{T_{f}}}^{i}=(\widehat{x}_{t+{T_{f}}}^{i},\widehat{y}_{t+{T_{f}}}^{i})$ (6)
where $\widehat{x}_{t+{T_{f}}}^{i}$ and $\widehat{y}_{t+{T_{f}}}^{i}$ represent the predicted horizontal and vertical coordinates of the $i$-th target vehicle at time step $(t+T_{f})$, respectively.
III MODEL ARCHITECTURE
This article introduces an Endpoint-based Planning-informed Network (EPN) for multimodal trajectory prediction by integrating vehicle driving status information with the future planning trajectory of the ego vehicle, the model architecture is shown in Figure 2. EPN consists of three main modules: the feature fusion encoding, the target’s endpoint prediction, and the LSTM decoder. The feature fusion encoding module separately encodes the information from the ego vehicle, target vehicle, and adjacent vehicles. The encoded features of the ego vehicle and adjacent vehicles are processed through a convolutional social pooling network to extract interaction information, generating social feature vectors. Simultaneously, the target vehicle’s encoded features are transformed into dynamic feature vectors via a fully connected layer. These vectors are concatenated to form environmental feature vectors. The target’s endpoint prediction module employs CVAE to predict potential target vehicle endpoints and refines them to improve accuracy. Finally, the LSTM decoder module combines the environmental and endpoint feature vectors to produce the target vehicle’s complete multimodal predicted trajectory.
Figure 2: Overall Architecture of EPN Model.
III-A Feature Fusion Encoding
The feature fusion encoding module is designed to output an environmental feature vector that integrates the spatiotemporal information of the driving scene. This module encodes not only the historical trajectories and states of all vehicles in the scene but also the planned trajectory of the ego vehicle, fully considering its potential influence on the future trajectory of the target vehicle. Given LSTM’s effectiveness in handling temporal data, we utilize LSTM networks to encode vehicle information. Due to the limitations of the dataset and challenges in obtaining complete driving status data, we focus on using each vehicle’s historical trajectory, speed, and acceleration as inputs. After processing this input information through an LSTM encoder, convolutional social pooling network, and fully connected layer, the final output is an environmental feature vector.
Since trajectory data, velocity, and acceleration have different dimensions, we apply distinct embedding layers with unique parameters for each, followed by separate LSTM encoders. Each vehicle’s historical trajectory and state data is initially passed through an embedding layer, then processed by an LSTM encoder, where the tensor from the final hidden layer represents the vehicle’s feature vector at the current time step. The specific computation process follows:
$h_{p}^{i}=LSTM(emb(x_{t}^{i},y_{t}^{i}))$ (7)
$h_{v}^{i}=LSTM(emb(v_{t}^{i}))$ (8)
$h_{a}^{i}=LSTM(emb(a_{t}^{i}))$ (9)
in equations (7) to (9), $t=\left({t-{T_{h}}+1,t-{T_{h}}+2,...,t}\right)$, $(x_{t}^{i},y_{t}^{i})$, $v_{t}^{i}$, and $a_{t}^{i}$ denote the position coordinates, velocity, and acceleration of the $i$-th vehicle at time step $t$, respectively. Here, $h_{p}^{t}$, $h_{v}^{t}$, and $h_{a}^{t}$ represent the feature vectors associated with the final hidden layer of the LSTM encoder for position, velocity, and acceleration. Once these three feature vectors are obtained, a fully connected layer calculates the intrinsic relationships among the vehicle’s position, velocity, and acceleration, integrating these features into a unified dimension. Additionally, a fully connected layer performs dimensionality reduction on the target vehicle’s feature vector, producing a dynamic feature vector that captures the target vehicle’s temporal characteristics. The following equation details this computation:
$e_{nei}^{t}=FC(h_{p}^{i},h_{v}^{i},h_{a}^{i}),i\in{A_{nei}}$ (10)
$e_{tar}^{t}=FC(h_{p}^{i},h_{v}^{i},h_{a}^{i}),i\in{A_{tar}}$ (11)
$e_{dyn}=FC(e_{tar}^{t})$ (12)
here, $e_{nei}^{t}$ represents the final encoded feature vector of the vehicles surrounding the target vehicle at time $t$, which includes encoded information about the ego vehicle, while $e_{tar}^{t}$ is the final encoded feature vector of the target vehicle at time $t$. $e_{dyn}$ denotes the dynamic feature vector of the target vehicle after dimensionality reduction.
To encode the planned trajectory of the ego vehicle, we use embedding layers and LSTM networks to process its planning information:
$e_{plan}^{t}=LSTM(emb(x_{t}^{i},y_{t}^{i}))$ (13)
To effectively model the interaction between the ego vehicle, adjacent vehicles, and the target vehicle—and to emphasize the influence of adjacent vehicles’ historical information and the ego vehicle’s planned future trajectory on the target vehicle’s future path—we introduce a convolutional social pooling network to simulate inter-vehicle interactions within the scene. First, a mesh centered on the target vehicle is established using a masking mechanism, which collects the feature vectors of surrounding vehicles into a social tensor. The masking calculation formula is as follows:
$mas{k_{i,j}}=\left\{\begin{array}[]{l}1,{\rm{}}if{\rm{}}gri{d_{i,j}}=1\\ 0,{\rm{}}if{\rm{}}gri{d_{i,j}}=0\end{array}\right.$ (14)
here $gri{d_{i,j}}$ indicates whether an adjacent vehicle occupies position $\left({i,j}\right)$ in the grid centered on the target vehicle. If $gri{d_{i,j}=1}$, it signifies the presence of an adjacent vehicle at that position, and $mas{k_{i,j}}$ is set to 1; otherwise, $mas{k_{i,j}}$ is set to 0. The resulting social tensor after masking adjacent vehicles is denoted as $tenso{r_{nei}}$, while the social tensor for the planned trajectory of the ego vehicle is represented as $tenso{r_{plan}}$.
TABLE I: Structure and Parameters of Convolutional Social Pooling Network
Layer Name Function & Parameter
Conv1 Conv2d(64,64,3)
Relu LeakyRelu(0.1)
MP1 MaxPool2d((3,3),2)
Conv2 Conv2d(64,16,(3,1))
MP2 MaxPool2d((2,2),(1,0))
Figure 3: The Structure of Convolutional Social Pooling Network.
The social tensor is subsequently input into the convolutional pooling network to extract interaction information between vehicles. Table I presents the structure and parameters of the convolutional social pooling network, which consists of two convolutional layers, two pooling layers, and a ReLU layer. Figure 3 illustrates the tensor transformation process within the convolutional social pooling network. In this architecture, each convolutional layer is followed by a ReLU layer, employing $LeakyRelu(0.1)$ as the activation function. The primary role of the convolutional layers is to extract interactive information from the social tensors through convolutional operations, while the pooling layers downsample features to reduce computational complexity and enhance the model’s translational invariance. For the social tensors representing adjacent vehicles and the ego vehicle planning, interaction information is extracted using the ”convolution-pooling-convolution” operation, resulting in the intermediate tensors $en{c_{nei}}$ and $en{c_{plan}}$. These tensors are then concatenated through a pooling layer to generate the social feature vector, denoted as $en{c_{social}}$, which encapsulates the spatial feature information of the target vehicle. The formulas are as follows:
$en{c_{nei}}=Relu(Conv2(MP1(Relu(Conv1(tenso{r_{nei}})))))$ (15)
$en{c_{plan}}=Relu(Conv2(MP1(Relu(Conv1(tenso{r_{plan}})))))$ (16)
$en{c_{social}}=MP2(concat(en{c_{nei}},en{c_{plan}}))$ (17)
Finally, we combine the dynamic feature vector of the target vehicle with the social feature vector to create the environmental feature vector, denoted as $enc$. The formula is as follows:
$enc=concat(en{c_{social}},en{c_{dyn}})$ (18)
This environmental feature vector encompasses all relevant information derived from the raw data. Subsequent modules will utilize this environmental feature vector to predict the potential endpoint and future trajectory of the target vehicle.
III-B Target’s Endpoint Prediction
After obtaining the environmental feature vectors, we employ a CVAE to predict the potential endpoint positions that the target vehicle may reach. The CVAE can generate more appropriate sampling points by constraining the input data information to the variational autoencoder. To enhance the accuracy of the endpoints generated by the CVAE, we incorporate the true endpoint coordinates of the target vehicle into the model input.
In this module, the model extracts endpoint features using different processes at various stages. During the training phase, distinct multilayer perceptrons (MLPs) are employed as the endpoint encoder, latent variable encoder, and latent variable decoder. First, the endpoint encoder encodes the target vehicle’s true endpoint. The resulting feature vector is concatenated with the environmental feature vector and passed to the latent variable encoder to produce a latent variable. The latent variable’s mean $\mu$ and standard deviation $\sigma$ are calculated, and Gaussian noise $z$ is sampled from a normal distribution $N(\mu,\sigma)$. Finally, $z$ is concatenated with the environmental feature vector and decoded by the latent variable decoder to generate the target vehicle’s endpoint. The formulas for the training phase are provided in equations (19) to (21):
$end\_feature={E_{end}}(G)$ (19)
$latent={E_{latent}}(concat(enc,end\_feature))$ (20)
$\overline{G}={D_{latent}}(concat(enc,z))$ (21)
where $G$ represents the true endpoint of the target vehicle, while ${E_{end}}$, ${E_{latent}}$, and ${D_{latent}}$ denote the endpoint encoder, latent variable encoder, and latent variable decoder, respectively. $end\_feature$ refers to the encoded endpoint feature vector, and $\overline{G}$ represents the predicted endpoint of the target vehicle. During the validation and testing phases, since the true endpoint of the target vehicle is unknown, we directly sample Gaussian noise $z$ from the latent space based on a normal distribution $N(0,{\sigma_{T}}I)$. The sampled noise $z$ is then concatenated with the environmental feature vector and decoded to generate potential endpoint predictions. Following the truncation trick in PECnet[36], we set the standard deviation ${\sigma_{T}=1.3}$. Additionally, by controlling the number of times the latent space is sampled, multiple possible endpoint coordinates for the target vehicle can be predicted.
Due to the significant error in directly predicting the target vehicle’s endpoint, we employ a endpoint correction mechanism to improve the accuracy of the prediction. After the first-stage prediction results are generated, the correction mechanism is applied to refine the predicted endpoint. Specifically, we first use a endpoint encoder to re-encode the predicted endpoint $\overline{G}$ of the target vehicle. The resulting feature vector is concatenated with the environmental feature vector and input into a endpoint decoder ${D_{end}}$, which has the same structure as the latent variable decoder ${D_{latent}}$. This decoder generates a deviation value that represents the $offset$ between the predicted endpoint and the true endpoint. Finally, the $offset$ is added to the first-stage predicted endpoint to obtain the corrected endpoint $\widehat{G}$. The detailed calculation process is shown in equations (22) to (24):
$refine\_end\_feature={E_{end}}(\overline{G})$ (22)
$offset={D_{end}}(concat(enc,refine\_end\_feature))$ (23)
$\widehat{G}=\overline{G}+offset$ (24)
III-C LSTM Decoder
In the LSTM decoder module, the endpoint encoder is used to encode the corrected endpoint $\widehat{G}$. The resulting encoded feature vector is then concatenated with the environmental feature vector and input into the LSTM decoder. This process generates complete trajectories for each corrected endpoint $\widehat{G}$, resulting in multi-modal trajectory prediction outcomes, the calculation process is shown in formulas (25) and (26):
$refined\_end\_feature={E_{end}}(\widehat{G})$ (25)
$\widehat{Y}=LSTM(concat(enc,refined\_end\_feature))$ (26)
Due to the significant variations in the absolute coordinate values of vehicles across different scenarios, we aim to mitigate the adverse effects of these absolute values on the performance of trajectory prediction models. To achieve this, we predict the relative displacement $\left({\delta x_{t+\tau}^{i},\delta y_{t+\tau}^{i}}\right)$ of the target vehicle over the next ${T_{f}}$ time steps, where $\tau\in\left({1,2,...,{T_{f}}}\right)$. By predicting the relative displacement, we can improve the accuracy of the predictions. Once the relative displacement is predicted, the absolute coordinates of the predicted trajectory can be calculated using the formulas (27) and (28):
$x_{t+T_{f}}^{i}=x_{t}^{i}+\sum\limits_{\tau=1}^{T_{f}}\delta x_{t+\tau}^{i}$ (27)
$y_{t+T_{f}}^{i}=y_{t}^{i}+\sum\limits_{\tau=1}^{T_{f}}\delta y_{t+\tau}^{i}$ (28)
here $x_{t}^{i}$ and $y_{t}^{i}$ represent the horizontal and vertical coordinates of the target vehicle at time step $t$, respectively.
III-D Loss Function
We use the trajectory prediction error ${L_{pred}}$ and the conditional variational autoencoder error ${L_{cvae}}$ as loss functions. The trajectory prediction error ${L_{pred}}$ is calculated as the mean squared error (MSE) between the predicted trajectory $\widehat{Y}$ and the true trajectory $Y$, as well as between the corrected predicted endpoint $\widehat{G}$ and the true endpoint $G$. The conditional variational autoencoder used for endpoint prediction is trained using Kullback-Leibler (KL) divergence as its loss function. The specific loss functions are defined in equations (29) and (30):
${L_{pred}}={L_{mse}}(Y,\widehat{Y})+{L_{mse}}(G,\widehat{G})$ (29)
${L_{cvae}}={D_{KL}}(N(\mu,\sigma)||N(0,1))$ (30)
IV EXPERIMENTS
IV-A Dataset and Data Preprocessing
We trained and evaluated our model on two publicly available, large-scale, real-world highway datasets: NGSIM[37] [38] and HighD[39].
(1) NGIMS dataset: The NGSIM dataset was collected through a project initiated by the U.S. Federal Highway Administration. It contains real highway vehicle trajectory data from US101 and I80, with each sub-dataset capturing 45 minutes of trajectory data in mild, moderate, and congested traffic conditions. The dataset records vehicle position, speed, acceleration, and type at a sampling frequency of 10 Hz.
(2) HighD dataset: The HighD dataset was recorded using drones equipped with cameras over six highways in Germany, providing an aerial perspective. This dataset benefits from advanced machine learning algorithms, ensuring high localization accuracy. It includes data from over 110000 vehicles, covering a total mileage of 44500 kilometers, with information on vehicle position, speed, and acceleration sampled at 25 Hz.
Given the different sampling frequencies and data characteristics of the two datasets, we performed data preprocessing to standardize them. For our experiments, we used a 3-second historical trajectory to predict a 5-second future trajectory, creating 8-second scene segments during preprocessing. To reduce data volume, the sampling frequency was downsampled to 5 Hz, yielding 40 time steps per scenario, with 15 steps for history (${T_{h}}=15$) and 25 steps for prediction (${T_{f}}=25$). The datasets were divided into training, validation, and testing sets in a 7:1:2 ratio.
IV-B Experimental Setup and Evaluation
The experimental setup for this study includes Python 3.7, PyTorch 1.7, and CUDA 11.7, with all experiments trained on an RTX 3080 GPU. The LSTM encoder is configured with a dimension of 64, while the LSTM decoder has a dimension of 128. The batch size is set to 64, and the model is trained for 15 epochs with a learning rate of 0.001. For comparison, the CL-LSTM and PiP methods used in this experiment categorize the driving intentions of the target vehicle into six classes, generating multimodal trajectory predictions based on these intentions. Accordingly, the number of predicted endpoints in the endpoint prediction module is set to $k=6$ in this experiment.
Root Mean Square Error (RMSE) is the primary evaluation metric used on the NGSIM and HighD datasets, thus, we employ RMSE for model comparison and ablation studies. Additionally, we use Average Displacement Error (ADE) and Final Displacement Error (FDE) to further evaluate the model’s predictive accuracy. These metrics provide a comprehensive assessment of predictive performance.
IV-C Model Comparison
We compare and evaluate EPN against the following trajectory prediction models:
(1) S-LSTM[18]: Uses an LSTM encoder-decoder for vehicle trajectory prediction, with a fully connected social pooling layer to predict future trajectories.
(2) CS-LSTM[17]: Builds on S-LSTM with a convolutional social pooling layer to capture interactions and incorporates multimodal trajectory prediction based on horizontal and vertical driving intentions.
(3) S-GAN[21]: Combines sequence prediction with GANs, generating multiple trajectory predictions and using the closest to the true future trajectory for evaluation.
(4) WSiP[22]: Inspired by wave superposition, this model aggregates local and global vehicle interactions for dynamic social pooling.
(5) PiP[29]: Considers ego vehicle planning’s impact on nearby vehicles’ trajectories using an LSTM encoder and a convolutional social pooling module.
(6) S-TF[32]: Utilizes a sparse Transformer for multimodal prediction, incorporating trajectory, velocity, and acceleration information along with driving intentions (left offset, right offset, or straight).
(7) EPN: Our proposed model, which predicts multimodal trajectories based on endpoint correction and ego vehicle planning, selecting the trajectory closest to the true future trajectory for evaluation.
Figure 4: RSME Bar Chart of Each Model on NGSIM Dataset.
Figure 5: RSME Bar Chart of Each Model on HighD Dataset.
TABLE II: RSME Comparison of Each Model on NGSIM Dataset
Prediction duration RSME
S-LSTM CS-LSTM S-GAN WSiP PiP S-TF EPN
1s 0.60 0.58 0.57 0.56 0.55 0.99 0.38
2s 1.28 1.26 1.32 1.23 1.18 1.43 0.79
3s 2.09 2.07 2.22 2.05 1.94 1.7 1.12
4s 3.10 3.09 3.26 3.08 2.88 2.02 1.56
5s 4.37 4.37 4.40 4.34 4.04 3.33 2.31
TABLE III: RSME Comparison of Each Model on HighD Dataset
Prediction duration RSME
S-LSTM CS-LSTM S-GAN WSiP PiP S-TF EPN
1s 0.19 0.19 0.30 0.20 0.17 0.75 0.08
2s 0.57 0.57 0.78 0.60 0.52 0.89 0.18
3s 1.18 1.16 1.46 1.21 1.05 1.05 0.32
4s 2.00 1.96 2.34 2.07 1.76 1.33 0.55
5s 3.02 2.96 3.41 3.14 2.63 1.75 0.91
Tables II and III present the RMSE values for different models on the NGSIM and HighD datasets, respectively. A lower RMSE indicates smaller prediction errors. Bold text highlights the method with the best performance among all compared models. As shown in Tables II and III, our proposed EPN model achieved the best predictive performance on both the NGSIM and HighD datasets.
When compared to models such as L-LSTM, CS-LSTM, S-GAN, and WSiP, which only consider historical trajectory information, the prediction accuracy of EPN is significantly improved. This demonstrates that incorporating the vehicle’s driving status information and self-planning plays a crucial role in trajectory prediction. Furthermore, compared to the PiP model, which also accounts for the impact of ego vehicle planning on trajectory prediction, EPN exhibits a smaller prediction error, highlighting the benefits of our enhanced endpoint prediction module in improving model performance.In comparison with the S-TF model, which exhibits the smallest prediction error among the competing methods, EPN reduced the average RMSE by 34.9% on the NGSIM dataset and 64.6% on the HighD dataset. Figures 4 and 5 provide bar charts that visually compare the RMSE evaluation metrics across various models for both the NGSIM and HighD datasets. These results clearly demonstrate that our proposed EPN model outperforms the other models by a significant margin.
In addition to comparing the predictive performance of various models using statistical metrics, we also used Average Displacement Error (ADE) and Final Displacement Error (FDE) to further evaluate the performance of EPN and other multimodal trajectory prediction methods based on driving intentions. The comparison methods are as follows:
(1) CS-LSTM (M): For the multimodal trajectory prediction results based on different driving intentions output by the CS-LSTM method, we select the trajectory closest to the true future trajectory and use it to calculate the evaluation metrics.
(2) PiP (M): Similarly, for the multimodal trajectory prediction results based on different driving intentions from the PiP method, we select the trajectory closest to the true future trajectory to calculate the evaluation metrics.
(3) EPN: Our proposed multimodal trajectory prediction method, based on endpoint correction and ego vehicle planning, selects the trajectory closest to the true future trajectory from the predicted results to calculate the evaluation metrics.
Figure 6: ADE/FDE Bar Chart of Each Model on NGSIM Dataset.
Figure 7: ADE/FDE Bar Chart of Each Model on HighD Dataset.
TABLE IV: ADE/FDE Comparison of Each Model on NGSIM Dataset and HighD Dataset
Prediction duration ADE/FDE
NGSIM HighD
CS-LSTM(M) PiP(M) EPN CS-LSTM(M) PiP(M) EPN
1s 0.19/0.35 0.19/0.34 0.12/0.24 0.09/0.16 0.08/0.14 0.04/0.07
2s 0.40/0.75 0.39/0.74 0.28/0.54 0.20/0.42 0.18/0.40 0.08/0.14
3s 0.58/1.13 0.58/1.11 0.41/0.76 0.34/0.79 0.37/0.81 0.12/0.24
4s 0.79/1.57 0.77/1.53 0.53/1.01 0.53/1.28 0.51/1.13 0.17/0.39
5s 1.03/2.41 1.00/2.21 0.69/1.58 0.75/1.88 0.69/1.64 0.24/0.63
Table IV compares the ADE and FDE of EPN with the multimodal trajectory prediction methods CS-LSTM (M) and PiP (M). Figures 6 and 7 present the ADE and FDE evaluation metrics in the form of bar charts for the NGSIM and HighD datasets.
The experimental results show that when ego vehicle planning information is incorporated, the ADE and FDE values of PiP (M) are lower than those of CS-LSTM (M). Compared to PiP (M), EPN achieves an average reduction of 30.7% in ADE and 30.4% in FDE on the NGSIM dataset, and a more substantial reduction of 64.5% in ADE and 64.3% in FDE on the HighD dataset. These results demonstrate that our proposed multimodal trajectory prediction method, based on endpoint correction, outperforms the method based on predicted driving intentions. The performance improvement is especially significant on the HighD dataset, where the positioning data is more accurate, leading to a more pronounced reduction in prediction error.
IV-D Ablation Experiment
To evaluate the impact of each module in the EPN model on its predictive performance, we conducted ablation experiments with the following specific experimental settings:
(1) PCS-LSTM: Adds ego vehicle planning information to the CS-LSTM model.
(2) PCS-LSTM (V): Builds on PCS-LSTM by incorporating the vehicle’s speed as a feature input.
(3) PCS-LSTM (V+A): Extends PCS-LSTM by encoding both the speed and acceleration of the vehicle as features, which are input into the model.
(4) PCS-LSTM (V+A+M): Uses the optimal trajectory from PCS-LSTM to calculate evaluation metrics.
(5) EPN (R): Adds a endpoint prediction module to the PCS-LSTM model, while removing the endpoint correction mechanism from the full model.
(6) EPN: The complete trajectory prediction framework proposed in this paper, based on endpoint correction and ego vehicle planning.
TABLE V: RSME/ADE/FDE Comparison of Ablation Experiments on NGSIM Dataset
Methods RMSE/ADE/FDE
1s 2s 3s 4s 5s
PCS-LSTM 0.55/0.20/0.37 1.19/0.43/0.87 1.95/0.69/1.44 2.90/0.98/2.13 4.07/1.30/2.98
PCS-LSTM(V) 0.45/0.14/0.29 1.08/0.36/0.77 1.82/0.60/1.32 2.74/0.88/1.99 3.87/1.20/2.82
PCS-LSTM(V+A) 0.42/0.13/0.26 1.03/0.33/0.73 1.74/0.57/1.27 2.62/0.84/1.92 3.72/1.15/2.72
PCS-LSTM(V+A+M) 0.40/0.12/0.25 0.94/0.31/0.66 1.46/0.50/1.04 1.88/0.70/1.45 2.70/0.92/2.10
EPN(R) 0.39/0.12/0.24 0.79/0.28/0.55 1.15/0.41/0.78 1.62/0.55/1.08 2.39/0.72/1.67
EPN 0.38/0.12/0.24 0.79/0.28/0.54 1.12/0.41/0.76 1.56/0.53/1.01 2.31/0.69/1.58
Table V presents the comparison results of RMSE, ADE, and FDE from the ablation experiments on the NGSIM dataset. Figure 8 provides a line graph that visually illustrates the impact of each module on the model’s predictive performance. Incorporating vehicle driving status information, such as speed and acceleration, improves trajectory prediction accuracy. Compared to the benchmark model, PCS-LSTM, the addition of speed as a feature in PCS-LSTM (V) results in average reductions of 6.6%, 11.7%, and 7.7% in RMSE, ADE, and FDE, respectively. Further including both speed and acceleration in PCS-LSTM (V+A) leads to average reductions of 10.6%, 16.1%, and 11.4% in these metrics. When comparing PCS-LSTM (V+A+M) to EPN (R), the latter shows an average reduction of 14.1%, 18.4%, and 21.5% in RMSE, ADE, and FDE, respectively. This highlights the significant impact of the endpoint prediction module on improving the performance of multimodal trajectory prediction. Finally, by comparing EPN (R) with the full EPN model, we observe that the addition of the endpoint correction mechanism contributes to further improvements in the model’s predictive performance.
(a) RSME line chart of ablation experiment in NGSIM dataset.
(b) ADE line chart of ablation experiment in NGSIM dataset.
(c) FDE line chart of ablation experiment in NGSIM dataset
Figure 8: Line graph of ablation experiment under NGSIM dataset.
Figure 9: Visualization Results in Left-turn, Straight-ahead and Right-turn Driving Scenarios.
IV-E Visualization
In addition to the quantitative analysis of the model’s trajectory prediction performance, we also conducted a qualitative analysis. Figure 9 presents the multimodal trajectory prediction results of EPN in left turn, straight, and right turn driving scenarios for the target vehicle, along with a comparison of the optimal predictions between the EPN and PiP methods. Different colored curves are used to represent various trajectories: the gray curve shows the historical trajectories of all vehicles, the red curve represents the ego vehicle’s future planned trajectory, the green curve represents the true future trajectory of the target vehicle, the purple curve shows the predicted trajectory of the PiP method, and the blue curve indicates the predicted trajectory of EPN.
From the visualizations, it is clear that the multimodal trajectory predictions of EPN closely align with the true future trajectory of the target vehicle, demonstrating strong prediction performance. While the predictions for straight driving are similar between EPN and PiP, in left and right turn scenarios, the EPN predictions more closely match the true trajectory than those of PiP. This is because, in turning scenarios, predicting the vehicle’s possible endpoint directly—rather than relying on ambiguous driving intentions—enables a more accurate estimate of the vehicle’s movement, resulting in predictions that better match the true trajectory.
V CONCLUSION
In this paper, we propose a trajectory prediction model based on ego vehicle planning. This model uses the historical trajectory, vehicle speed, acceleration, and planned future trajectory of the ego vehicle as inputs, and outputs a multimodal prediction of the target vehicle’s future trajectory. By incorporating the planned trajectory of the ego vehicle, the interaction between the ego vehicle’s plan and the predicted trajectory of the target vehicle can be simulated, thereby more realistically replicating the driving scenario and improving the accuracy of trajectory prediction for the target vehicle. We introduce a target’s endpoint prediction module based on a CVAE to address the issues of intention ambiguity and large prediction errors typically found in trajectory prediction methods based on driving intentions. This module first predicts multiple potential endpoints for the target vehicle, and then refines the accuracy of these predictions through a endpoint correction mechanism. The module subsequently generates a complete multimodal trajectory based on the predicted endpoints. Experimental results demonstrate that our method achieves superior trajectory prediction accuracy compared to methods based solely on driving intentions.